OpenStudy (watchmath):

Calculate \[\sum_{n=1}^\infty \frac{n-1}{n!}\]

OpenStudy (anonymous):

0

OpenStudy (anonymous):

it has to be positive though

OpenStudy (watchmath):

Can't be zero since it is a positive series starting from \(n=2\)

OpenStudy (anonymous):

I can't prove it, but the sum seems to approach one, in that every successive sum is (n!-1)/n!

OpenStudy (watchmath):

Well observation is half of the proof thebestpig :). You just need to push a little bit more :D.

OpenStudy (anonymous):

gimme a second - i'll do an induction thing

OpenStudy (anonymous):

the base step is \[(1!-1)/1! = 0\] which is in fact the first term, and \[\frac{(n!-1)}{n!}+\frac{n}{(n+1)!}=\frac{(n!-1)(n+1)+n}{(n+1)!}=\frac{(n+1)!-n-1+n}{(n+1)!}\] \[=\frac{(n+1)!-1}{(n+1)!}\] and it looks like we're done.

OpenStudy (anonymous):

oh wait one more thing

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \frac{n!-1}{n!}=1\]

OpenStudy (watchmath):

Awesome! :)

OpenStudy (watchmath):

We can also using telescoping sum. Note that \[\frac{n-1}{n!}=\frac{1}{(n-1)!}-\frac{1}{n!}\]

OpenStudy (anonymous):

ha, /facepalm/

OpenStudy (anonymous):

well, I enjoy using induction :)

OpenStudy (anonymous):

Here's another way: a little relabeling gives (check it!) \[ \sum_{n=1}^\infty \frac{n-1}{n!} = \sum_{n=0}^{\infty}\frac{1}{(n+2)n!} \] Now \[ e^x = \sum_{n=0}^\infty \frac{x^n}{n!}, \] so \[ xe^x = \sum_{n=0}^\infty \frac{x^{n+1}}{n!}, \] so \[ \int xe^x = xe^x-e^x +C = \sum_{n=0}^\infty \frac{x^{n+2}}{(n+2)n!} \] Plug in \(x=0\) to find that \(C =1\), so \[ 1 e^1 -e^1 +1 = 1 = \sum_{n=0}^\infty \frac{1}{(n+2)n!}. \]